West Palm Beach hosted a local middle school basketball tournament where each
team played three games. In the last game, the Tigers scored 56 points, which
was 17 points different from what they scored in the second game. The score of
the second game was 12 points different from the score of the first game. What
could have been the amount of points the Tigers scored in the first game?

Answer by ikleyn(52790)

.
West Palm Beach hosted a local middle school basketball tournament where each
team played three games. In the last game, the Tigers scored 56 points, which
was 17 points different from what they scored in the second game. The score of
the second game was 12 points different from the score of the first game. What
could have been the amount of points the Tigers scored in the first game?
~~~~~~~~~~~~~~~~~~~

The third game score 56 points produces two possible scores for the second game

    56+17 = 73  and  56-17 = 39.



The score 73 of the second game produces two possible scores for the first game

    73+12 = 85  and 73-12 = 61.



The score 39 of the second game produces two possible scores for the first game

    39+12 = 51  and 39-12 = 27.


So, the ANSWER is these four possible scores for the first game: 27, 51, 61 and 85.

Solved.